Please help me understand how to work this problem.
Evaluate each function at the given values of the independent variable and simplify.
f(x) = 4x + 5
a. f(6) b. f(x+1) c. f(-x)

Answer by jessica43(140)

For these type of problems, you need to plug the value of x into the equation.
a)For this part, the problem is asking you what does the function/equation equal when x=6. So plug 6 in for x:
f(x) = 4x +5
f(6) = 4(6) +5
f(6) = 24 + 5
f(6) = 29
b) Now plug in (x+1) for x:
f(x) = 4x + 5
f(x+1) = 4(x+1) + 5
f(x+1) = 4x + 4 + 5
f(x+1) = 4x + 9
c)Now plug in -x for x:
f(x) = 4x + 5
f(-x) = 4(-x) + 5
f(-x) = -4x + 5